OpenStudy (anonymous):

Find the cosine of both angle A and angle B. A. cos A = 5/13; cos B = 12/13 B. cos A = 12/13; cos B = 5/13 C. cos A = 5/12; cos B = 12/5 D. cos A = 12/5; cos B = 5/12

OpenStudy (anonymous):

@JakeV8

OpenStudy (anonymous):

may I answer this?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

cos = adjacent side / hypotenuse cos A=24/10=12/5 cos B=10/26=5/13

OpenStudy (anonymous):

do i solve those ?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

wait is that supposed to be D ? the last one is supposed to say 5/12

OpenStudy (anonymous):

cos A should be 24/26 =12/13 I did it wrong

OpenStudy (anonymous):

sorry for that

OpenStudy (anonymous):

Those aren't quite correct... There you go... good catch :)

OpenStudy (anonymous):

so my answer would be B then right ?

OpenStudy (anonymous):

yes. you got it!

OpenStudy (anonymous):

thanks (:

OpenStudy (anonymous):

Wait...isn't that the cosecant?